# Silent Guardian: Protecting Text from Malicious Exploitation by Large Language Models

### v4

@@ Abstract
The rapid development of large language models (LLMs) has yielded impressive success in various downstream tasks. However, the vast potential and remarkable capabilities of LLMs also raise new security and privacy concerns if they are exploited for nefarious purposes due to their open-endedness. For example, LLMs may be used to plagiarize or imitate writing, thereby infringing the copyright of the original content, or to create indiscriminate fake information based on a certain source text. In some cases, LLMs can even analyze text from the Internet to infer personal privacy. Unfortunately, previous text protection research could not foresee the emergence of powerful LLMs, rendering it no longer effective in this new context.
@@
The comprehensive experimental evaluation demonstrates that SG can effectively protect the target text under various configurations and achieve almost 100% protection success rate in some cases. Notably, SG also exhibits relatively good transferability and robustness, making its application in practical scenarios possible.
@@ I Introduction
Recent advances in large language models (LLMs) have led to impressive performance in a variety of downstream language tasks [1], such as holding natural conversation [2], text and code generation [3, 4], and reading comprehension [5]. By automating tedious tasks and readily providing comprehensive information, they are expected to increase the productivity of society dramatically. However, while bringing convenience to people, this powerful ability also creates new security and privacy risks. For example, malicious users can use LLMs to exploit internet text to engage in illegal activities. Recent research [6] has indicated that by leveraging individual statements from social media, LLMs such as GPT-4 can accurately infer personal information such as gender, income, and location. This exposes personal privacy to enormous risks. In addition, the emergence of LLMs has automated the process of plagiarists and specialized artificial intelligence article rotation tools already exist [7], which poses a major challenge to copyright protection. Additionally, the capability of LLMs to mass-produce targeted rumors based on source texts [8] also makes governance in cyberspace increasingly difficult. It is worth noting that when LLMs have the ability to actively retrieve Internet texts, e.g., Bing Chat, highly automated malicious behaviors under the instructions of malicious users become possible, and the above risks are further amplified.
@@ V-B3 The Relationship between the Time and Hyperparameters for Constructing a Single TPE
Figure 6: The heatmap of the relationship between the replacement set size, batch size, and construction time. The masked areas indicate scenarios where, even after 100 epochs, the PSR did not reach 0.9, and the unit of the numbers in the figure is minutes.
@@ V-B5 Why Truncation Works
Figure 8: The optimization result of six text examples targeting different tokens.
@@ VI-C2 Discussion on the Robustness of TPE
The results indicate that the effectiveness of TPE protection deteriorates as the perturbation ratio increases. However, excessively high ratios of perturbation can cause significant loss to the original text.

### v6

@@ Abstract
The rapid development of large language models (LLMs) has yielded impressive success in various downstream tasks. However, the vast potential and remarkable capabilities of LLMs also raise new security and privacy concerns if they are exploited for nefarious purposes due to their open-endedness. For example, LLMs may be used to plagiarize or imitate writing, thereby infringing the copyright of the original content or to create indiscriminate fake information based on a certain source text. In some cases, LLMs can even analyze text from the Internet to infer personal privacy. Unfortunately, previous text protection research could not foresee the emergence of powerful LLMs, rendering it no longer effective in this new context.
@@
The comprehensive experimental evaluation demonstrates that SG can effectively protect the target text under various configurations and achieve almost 100% protection success rate in some cases. Notably, SG also exhibits relatively good transferability and robustness, making its application in practical scenarios possible. Our code is available at https://github.com/weiyezhimeng/Silent-Guardian.
@@ I Introduction
Recent advances in large language models (LLMs) have led to impressive performance in a variety of downstream language tasks [1], such as holding natural conversation [2], text and code generation [3, 4], log anomaly detection [5] and reading comprehension [6]. By automating tedious tasks and readily providing comprehensive information, they are expected to increase the productivity of society dramatically. However, while bringing convenience to people, this powerful ability also creates new security and privacy risks. For example, malicious users can use LLMs to exploit internet text to engage in illegal activities. Recent research [7] has indicated that by leveraging individual statements from social media, LLMs such as GPT-4 can accurately infer personal information such as gender, income, and location. This exposes personal privacy to enormous risks. In addition, the emergence of LLMs has automated the process of plagiarists and specialized artificial intelligence article rotation tools already exist [8], which poses a major challenge to copyright protection. Additionally, the capability of LLMs to mass-produce targeted rumors based on source texts [9] also makes governance in cyberspace increasingly difficult. It is worth noting that when LLMs have the ability to actively retrieve Internet texts, e.g., Bing Chat, highly automated malicious behaviors under the instructions of malicious users become possible, and the above risks are further amplified.
@@ V-B3 The Relationship between the Time and Hyperparameters for Constructing a Single TPE
Figure 6: The heatmap of the relationship between the replacement set size, batch size, and construction time. The masked areas indicate scenarios where, even after 100 epochs, the PSR did not reach 0.9, and the unit of the numbers in the figure in minutes.
@@ V-B5 Why Truncation Works
Figure 8: The optimization result of six text examples targeting different tokens. Each line plot corresponds to the optimization results for a specific set of tokens targeted in the text. The horizontal axis Epoch represents the number of optimization rounds, and the vertical axis SR (success rate) represents the probability of the model outputting the token, calculated similarly to PSR. The legend above each plot shows, from left to right, the top five or six tokens with the highest SR rankings before optimization.
@@ VI-C2 Discussion on the Robustness of TPE
The results indicate that the effectiveness of TPE protection deteriorates as the perturbation ratio increases. However, excessively high ratios of perturbation can cause significant information loss to the original text.
@@ VI-C6 Other Truncation Method
VI-D Evasion Attack
In this section, we will discuss the impact of three potential evasion attacks on TPE.
VI-D1 Adversarial Prompt Engineering
Against the TPE, an attacker can construct adversarial prompts to make the LLM regain the ability to output text. For example, an attacker can add text before and after the TPE to disrupt the TPE performance.
Effective adversarial prompts can be divided into two categories, one usually requiring more complex manual construction [48], while the other uses information such as gradients to optimize the construction process [19, 20, 21, 23, 22].
For the first type of attack, the involvement of human labor makes the cost of the attack rise. If only pre-designed text is used to reach the automation goal, the attack will be less effective. The experiment in Section V-B1 is similar to this scenario, which shows that the effect of adding prompts before and after the text on the performance of TPE decreases as the length of the TPE increases, and in addition, the addition of extraneous prompts may affect the model performance.
For the second type of attack, the attacker is able to construct adversarial prefixes and suffixes through gradient information, which in turn reduces the probability of sampling the end token in the first round. However, since lowering the probability of the end token does not correspond to the actual semantics, this approach may cause the LLM responses to deviate from the original question. In addition, this type of approach usually requires a large time overhead.
VI-D2 Simple Text Perturbation
Against the TPE, an attacker is able to degrade the protective effect of TPE through simple text perturbations. Examples include randomly adding words to the text, randomly deleting words from the text, randomly replacing words in the text, or segmenting the text. The experiment in Section VI-C2 is similar to this scenario, and the results show that increasing the text perturbation ratio can degrade the effectiveness of TPE protection, but it also causes information loss of original text, such as structural information, semantic information of the text.
VI-D3 Retokenization
Retokenization refers to encoding the inputs to the model in different tokenization methods [49]. In the simplest case, we break tokens apart and represent them using multiple smaller tokens. For example, the token “studying” has a broken-token representation “study”+“ing”, among other possibilities. Through the method of retokenization, an attacker is able to reduce the adversarial nature of the TPE, thus achieving the goal of reducing its performance. However, retokenization will degrade the performance of the model to some extent.